OpenStudy (dcher03):

Please help, I need an explanation on how to do this I will medal best answer. Winton wants to find the volume of a cube. He uses the formula for volume of a cube V=S 3, where S is the length of one side of the cube. If S =0.5 centimeter, what is the volume of the cube?

OpenStudy (mww):

this is just substitution. The formula says \[v = s^3\] which simply means the volume of the cube is found by taking the cubed number of the side. So do this.

OpenStudy (dcher03):

so, v= 0.5 times 3?

OpenStudy (mww):

no. do you know what a cubed number is? or what a square number is?

OpenStudy (ask_if_im_a_ghost):

S^3 means to multiply "S" 3 times. So V=S^3=S*S*S=(0.5 cm)*(0.5 cm)*(0.5 cm)=0.125 cm^3

OpenStudy (dcher03):

oh, sorry I don,t know what I was thinking, I am supposed to do 0.5 times 0.5 times 0.5

OpenStudy (mww):

yeah cubed means you multiply the number by itself 3 times, rather than times the number by three (which is really s + s + s)

OpenStudy (dcher03):

thnx. : )